OpenStudy (anonymous):

George has an annuity that pays $1,250 at the beginning of each quarter. If the economy grows at a rate of 3.49% annually, what is the value of the annuity if he received it in a lump sum now rather than over a period of nine years?

OpenStudy (anonymous):

@phi I am not sure which formula to use. I think I would do PV because it says if he received a lump sum right?

OpenStudy (phi):

This looks like the problem that confused me before. http://openstudy.com/users/jessica52595#/updates/50fff5d1e4b0426c6368597f The Present Value Annuity Due formula \[ PV= pymt \left(\frac{1}{i} - \frac{1}{i(1+i)^{nt-1} }+1\right ) \] I got it to work (match up with http://www.bankrate.com/calculators/investing/annuity-calculator.aspx if I subtract 1 from n*t Here payment is 1250 i = 0.0349/4 n=4, t=9 nt-1 is 35

OpenStudy (phi):

I got 38811.09